OpenStudy (anonymous):

Does the table represent a linear or an exponential function? Explain. x=1 2 3 4 y=16 64 256 1024

OpenStudy (anonymous):

at a glance the y values are increasing at a much greater rate than the x values. this is typical of a exponential function

OpenStudy (anonymous):

If the function was linear, the change in Y would always be the same. For example Y would increase or decrease by the same value for ever corresponding change in X by 1. So a linear function would be: X: 1,2,3,4 Y: 16,32,48,64. Y increase by 16 each time X increases by 1. This is an exponential function, we can see this because \[4^2=16\] \[4^3=64\] \[4^4=256\] and \[4^5=1024\]. Notice that the variable is in the exponent, each time the exponent is increasing by 1 not the base. This is what defines an exponential function. 4^x. 4 as a base remains the same, x is what changes; the exponent is X.